A man can row a boat with 4kmphin still water.if he is crossing a river where the current is 8kmph .the time he will cross?(width of river 8km).the river in shortest path is n/√3 hour, where n=?

Grade:12th pass

1 Answers

Krishna Kumar manchaala
5 years ago
Please refer the below problem which is same as it.....

Speed of the boat in still water = 5 km /hr

Width of the river = 1 km

Time taken by the boat to cross the river = 15 minutes = 15/60 hr = 1/4 hr

Let the velocity of the river be x km/hr.

Distance covered by the boat in 1/4 hr = 5/4 km = 1.25 km

Because of the flow of the river the boat will move in the direction of AC, which is the shortest possible path for the boat.

∴ AC = 1.25 km

AB (Width of the river) = 1 km

BC is the other bank of the river so the width AB of the river will be perpendicular to BC .i.e., ∠B = 90°

In ΔABC by Pythagoras theorem,

(AB)2 + (BC)= (AC)2

⇒ (1)2 + (BC)2 = (1. 25)2

⇒ (BC)2 = (1.25)2 – (1)2

⇒ (BC)2 = 1.5625 – 1

⇒ (BC)2 = 0.5625

⇒ BC = 0.75 km

Which means that distance covered by the river water in 15 minutes is 0.25 km.

∴ Velocity of the river = 0.75/(15/60) km/hr

 = 3 km/hr